OpenStudy (anonymous):

Find critical values at g(θ) = 36θ − 9 tan θ

OpenStudy (goformit100):

use allide angles

OpenStudy (anonymous):

take the first derivative \[g'(\theta)=36-(\sec \theta)^{2}\]then factor and set equal to zero\[9(4-(\sec \theta)^{2})=0\]9=0 doesn't work, so \[4-\sec ^{2}\theta=0\]solve for theta